OpenStudy (madinaxxx):

The CO2 produced, collected and cooled to O⁰C at 1atm. It is found to take up approximately 22.4 L. How many molecules of CO2 were produced?

OpenStudy (madinaxxx):

@ParthKohli

OpenStudy (abb0t):

For this, assume ideal behavior for the molecules. So, in the problem, you are given: \(\sf \color{red}{Tempetaure ~[T] = 273.15~K}\) \(\sf \color{blue}{Pressure ~[P] = 1~atm}\) \(\sf \color{green}{Volume ~[V]= 22.4~L}\) So, you're going to use: \(\sf \color{purple}{PV=nRT}\) Solve for \(n\), which is moles (this is algebra, rearrange the formula to solve for what you want). Also, keep in mind you're gas constant, \(\sf \color{orange}{8.314 \frac{L ~atm}{mol~K}}\) Once you have moles, use avogadro's number: \(\sf \color{green}{6.022 \times 10^{23} \frac{molecules}{mole}}\) to find the amount there are.

OpenStudy (madinaxxx):

I'm sorry I don't understand. Can you explain further please? @abb0t

OpenStudy (abb0t):

What don't you understand, where are you lost? I explained this quite thoroughly.

OpenStudy (madinaxxx):

@Preetha